OpenStudy (anonymous):

For 2x – 3y = –7, when x = 4, What is the value of y?

OpenStudy (ivettef365):

substitute the x for the 4, so you have 2(4) -3y = -7, now solve for y

OpenStudy (anonymous):

an what would you get

OpenStudy (ivettef365):

do you know how to solve it??

OpenStudy (anonymous):

yess for my answer I got 5

OpenStudy (anonymous):

i think i did something wrong

OpenStudy (ivettef365):

ok lets check 8 - 3y = -7 -8 -8 ------------- -3y = -15

OpenStudy (ivettef365):

-3y = -15 --- --- = 5 -3 -3

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

so i had it right...yaaaaayyyyy...lol thank you thank you

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

2x - 3y = -7 when x = 4 2(4) - 3y = -7 8 - 3y = -7 -8 -8 --- ---- -3y = -15 --- ---- -3 -3 y = ?